OpenStudy (anonymous):

what is the degree of the monomial 6x7y5

OpenStudy (anonymous):

12

OpenStudy (anonymous):

explain to me how is that the answer

OpenStudy (anonymous):

Degree of the given polynomial = degree of x + degree of y = 7+5 = 12

OpenStudy (anonymous):

if we consider all the variables to be the same .... say a 6 a7 a5, the a7 * a5 = a(7+5) just a good trick of multivariables confuse you